1.1) y^-2 dy/dx + y^-1 = 2x

1.2) (y^2e^xy^2 +4x^3)dx + (2xye^xy^2 - 3y^2)dy = 0
1
Expert's answer
2020-03-09T13:59:54-0400

"1.1) \ \frac{1}{y^2} \frac{dy}{dx}+\frac{1}{y}=2x"

"\ -\frac{1}{y^2} \frac{dy}{dx}-\frac{1}{y}=-2x"


"v(x)=\frac{1}{y(x)}, \ \frac{dv}{dx}=-\frac{1}{y^2(x)}\frac{dy}{dx}"


Substitute "v(x)" :

"\frac{dv }{dx}-v=-2x"


Multiply by "e^{-x}" both sides of equation:

"e^{-x}\frac{dv}{dx} -e^{-x}v=-2xe^{-x}"

"e^{-x}\frac{dv}{dx} +v\frac{d}{dx}(e^{-x})=-2xe^{-x}"

"\frac{d}{dx}(ve^{-x})=-2xe^{-x}"

"ve^{-x}=\int (-2xe^{-x})dx=2xe^{-x}+2e^{-x}+C"


Return to "y(x):"

"\frac{1}{y}e^{-x}= 2xe^{-x}+2e^{-x}+C"

"y=\frac{e^{-x}}{2xe^{-x}+2e^{-x}+C}=\frac{1}{2x+2+Ce^{x}}"


Answer: "y(x)=\frac{1}{Ce^x+2(x+1)}"


"1.2) (y^2e^{xy^2}+4x^3)dx+(2xye^{xy^2}-3y^2)dy=0"

"P(x,y)=y^2e^{xy^2}+4x^3, \ \ Q(x,y)=2xye^{xy^2}-3y^2"

"\frac{\partial P(x,y)}{\partial y} =\frac{\partial Q(x,y)}{\partial x}=2ye^{xy^2}+2xy^3e^{xy^2}"

So, we will find function "f(x,y):"

"\frac{\partial f(x,y)}{\partial x}=y^2e^{xy^2}+4x^3, \ \ \frac{\partial f(x,y) }{\partial y}=2xye^{xy^2}-3y^2"

"f(x,y)=\int (y^2e^{xy^2}+4x^3)dx= e^{xy^2}+x^4+\phi (y)"

"f(x,y)=\int (2xye^{xy^2}-3y^2)dy =e^{xy^2} -y^3+\psi (x)"


"f(x,y)=e^{xy^2}+x^4-y^3"

"\frac{\partial f(x,y)}{\partial x} dx+\frac{\partial f(x,y)}{\partial y} dy=0 \ \Rightarrow f(x,y)\equiv C"


Answer: "e^{xy^2(x)}+x^4-y^3(x)=C"